OpenStudy (anonymous):

What constant term should be used to complete the square? x^2 - 7x + _____ = 4

OpenStudy (btaylor):

(x+a)^2 = x^2 + 2ax + a^2. So a = -3.5, and -3.5^2 = 12.25